Could you please show me the step by step solution to this problem? Thank you
2(m+4)+3=3(m+2)+1m=?

Answer by checkley75(3666)

2(M+4)+3=3(M+2)+M
2M+8+3=3M+6+M
2M-3M-M=6-11
-2M=-5
M=-5/-2
M=2.5 ANSWER.
PROOF
2(2.5+4)+3=3(2.5+2)+2.5
2*6.5+3=3*4.5+2.5
13+3=13.5+2.5
16=16